If fixed costs are $320,000, the unit selling price is $72, and the unit variable costs...

  1. If fixed costs are $320,000, the unit selling price is $72, and the unit variable costs are $50, what are the old and new break-even sales in units (rounded to a whole number) if the unit selling price increases by $4?

    a.14,545 units and 4,444 units

    b.4,444 units and 14,545 units

    c.6,400 units and 11,808 units

    d.14,545 units and 12,308 units
Answer #1

Answer: Option ( d ) 14,545 units and 12,308 units

Explanation:

Break-even-point (Units) + Contribution margin per unit = Fixed cost 320000 320000 Old New 22 Break even point 14545 12308

Old : Contribution margin per unit = 72 - 50 = 22

When sales rice increased by $4

New : Contribution margin per unit = $76 - 50 = 26